OpenStudy (anonymous):

(^3sqrt24)/(^3sqrt15)

OpenStudy (anonymous):

\[\sqrt[3]{24}\div \sqrt[3]{15}\]

jhonyy9 (jhonyy9):

cuberoot(24/15) = cuberoot(8/5)=2/cuberoot5 = (2cuberoot5)/5

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

\[\frac{\sqrt[3]{24}}{\sqrt[3]{15}}=\frac{2}{5^{1/3}}=1.16961 \]------------------ \[\frac{2 \sqrt[3]{5}}{5}=\frac{2}{5^{2/3}}=0.68399 \]